240 dm3 of air at s.t.p is compressed to 17 atmospheres in a compression tube. Determine the volume in dm3 occupied in the tube.


1
Expert's answer
2022-02-28T11:01:46-0500

p1 * V1 = p2 * V2

So, V2 = (p1 * V1) / p2 = ( 1atm * 240dm3) / 17atm = 14.12 dm3

Amswer: 14.12 dm3